2. The equilibrium constant for the production of ammonia, used as fertilizer, is 3.65x 108 at 25 ⁰C. N2(g) + 3 H2(g) ⇋ 2 NH3(g) . Calculate the equilibrium concentration of hydrogen gas, if at equilibrium the concentrations of nitrogen and ammonia are 2.56 M and 5.76 M 3 The equilibrium constant of oxalic acid present in many fruits at 25 ⁰C. is 6.5 x 10-2. The reaction is C2H2O4(aq) ⇋ H + (aq) + C2HO4 - (aq) What is the concentration of H+ at 25⁰C if the concentration of C2H2O4 and C2HO4 - at equilibrium are 3.0111 M and 0.44415 M respectively? 


1
Expert's answer
2021-12-03T12:51:02-0500

The equilibrium constant for the production of ammonia, used as fertilizer, is 3.65x 108 at 25 ⁰C. N2(g) + 3 H2(g) ⇋ 2 NH3(g) . Calculate the equilibrium concentration of hydrogen gas, if at equilibrium the concentrations of nitrogen and ammonia are 2.56 M and 5.76 M

3.65x 108=394.2

Nitrogen=2.56M

Ammonia=5.76M

Total concentration=2.56M+5.76M

=8.32


The equilibrium constant of oxalic acid present in many fruits at 25 ⁰C. is 6.5 x 10-2. The reaction is C2H2O4(aq) ⇋ H + (aq) + C2HO4 - (aq) What is the concentration of H+ at 25⁰C if the concentration of C2H2O4 and C2HO4 - at equilibrium are 3.0111 M and 0.44415 M respectively?

C2H2O4=3.0111M

C2HO4=0.44415 M

6.5 x 10-2=0.0065

Total Equilibrium=3.0111+0.44415=3.45524

3.45524/0.0065=531.58